# AI/ML based Joint Source and Channel Coding for HARQ-ACK Payload

### v1

@@ XI Acknowledgments
The authors would like to thank Peter Gaal, Mike McCloud, Brian Banister and Hari Sankar for their valuable inputs and constructive feedback over the course of this project.
A Bitwise UEP for 1 bit
Consider the transmission of 1 bit, such that ACK is transmitted as $\sqrt{P_{1}}$ and NACK as $\sqrt{P_{0}}$. Since $k=1$, we can rewrite the decoding rule in (14) as
$\frac{\|y+\sqrt{P_{0}}\|^{2}}{2\sigma^{2}}\geq\ln\bigg(\frac{\beta}{1-\beta}\bigg)-\frac{\|y-\sqrt{P_{1}}\|^{2}}{2\sigma^{2}}.$ (48)
Observe that the $\|y\|^{2}$ cancels out, so that we can rewrite (48) as
$y\leq\frac{\sqrt{P_{1}}-\sqrt{P_{0}}}{2}+\frac{\sigma^{2}}{\sqrt{P_{1}}-\sqrt{P_{0}}}\ln\bigg(\frac{\beta}{1-\beta}\bigg).$ (49)
Hence, in the 1-bit case, the Bitwise UEP decoder simply reduces to comparing $y$ with a (SNR dependent) threshold $\gamma(\mathrm{SNR})$. Given a target NACK error rate $P(\mathrm{ACK}|\mathrm{NACK})=Q(\gamma(\mathrm{SNR})+\sqrt{P_{0}/\sigma^{2}})=\delta_{0}$, the entire RHS in (49) must equal $Q^{-1}(\delta_{0})-\sqrt{P_{0}/\sigma^{2}}$, and hence $\beta$ can be computed as a function of SNR and $\delta_{0}$.
@@ B Learnt PS - Training and Results
We then replace $\sqrt{\alpha_{m}}\mathbf{c}_{m}$ in Algorithm 1 by $\sqrt{\kappa_{\mathbf{\Pi}}}\mathbf{c}_{u,m}$, and the remainder of the training process remains unchanged. If we set $N=1$ and train on $(p,\rho)=(0.9,0)$, we will converge to the Learnt PS depicted in Fig. 7. If we perform free-lunch training as desrcibed in Section VII-D, the PS learnt (labeled Robust PS) is very similar to Arithmetic PS and is depicted in Fig. 13 for $p=0.9$ and $\rho=\{0,0.5,0.9\}$.
@@
C Impact of varying $\beta$ on $\mathrm{SNR}_{\mathrm{th}}$
In Section V-A, we presented an empirical rule for setting $\beta$ as per (15). Fig. 14 illustrates the impact of varying $\beta$. If $\beta=0.5\implies\beta/(1-\beta)=1$, then both ACK and NACK are equally protected, hence their BER curves coincide in Fig. 14a. As we lower $\beta$ to 0.25, NACK begins to achieve a lower BER than ACK in Fig. 14b. However, we observe that the SNR required to achieve a 0.1% NACK BER is higher than that required to achieve 1% ACK BER. Finally, at $\beta=1/11$ in Fig. 14c, we observe that the two aforementioned SNRs coincide, which implies that this value of $\beta$ minimizes $\mathrm{SNR}_{\mathrm{th}}$. Furthermore, we observe that $\mathrm{BER}_{\mathrm{NACK}}/\mathrm{BER}_{\mathrm{ACK}}=0.1$ for a band of SNRs around $\mathrm{SNR}_{\mathrm{th}}$. Similarly, we observe that the BER ratio is $1/3$ and 1 in Fig. 14b and 14a respectively. We observed this to be true for other values of $p$ as well, hence we arrived at the rule in (15).
@@ C Impact of varying $\beta$ on $\mathrm{SNR}_{\mathrm{th}}$
D Derivation of (25)
Note that $\mathbf{\Sigma}_{11}=\epsilon^{2}g^{2}\mathbb{E}[\mathbf{h}_{d,l}\mathbf{h}^{H}_{d,l}]$ ($l$ is the receiver antenna index), from which the formula for $\mathbf{\Sigma}_{11}$ in (25) follows. To compute $\mathbf{\Sigma}_{12}$, we have
$\mathbf{\Sigma}_{12}$ $=\epsilon g\mathbb{E}[\mathbf{h}_{d,l}\mathbf{y}^{H}_{p,l}]$ (51)
$=\epsilon g\mathbb{E}[\mathbf{h}_{d,l}\epsilon(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})^{H}]$
$=\epsilon^{2}g^{2}\sqrt{\beta_{m}}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{p}},$
@@ D Derivation of (25)
$=\mathbb{E}[\epsilon(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})\epsilon(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})^{H}]$
$=\epsilon^{2}g^{2}\beta_{m}\mathbf{R}_{\mathbf{h}_{p}\mathbf{h}_{p}}+\epsilon^{2}\sigma^{2}\mathbf{I}_{N_{p}}.$
E Derivation of (36)
@@
F Impact of varying ChEST SNR on RMMSE ChEST

### v2

@@ XI Acknowledgments
The authors would like to thank Yi Huang, Peter Gaal, Mike McCloud, Brian Banister and Hari Sankar for their valuable inputs and constructive feedback over the course of this project. We would also like to sincerely thank the reviewer who took the effort to provide in entirety the correct derivation for Appendix C.
A Group Permutation Symmetry
Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$ be a collection of input sequences, and let $\pi$ be any permutation of the indices $\{1,\ldots,n\}$. A function $f:\mathcal{X}^{n}\rightarrow\mathcal{Y}^{n}$ is said to be permutation equivariant if, for every permutation $\pi$,
$f(\pi\mathbf{X})=\pi f(\mathbf{X}).$ (47)
For example, consider the transformer architecture depicted in Fig. 2. It has been shown that transformers without positional embedding are permutation equivariant w.r.t token permutations (refer Theorem 4.1 in [xu2024permutation]). This implies that given a set of $A$ input tokens $\mathbf{x}$ and the transformer output for this sequence $H(\mathbf{x})$, the transformer output for any token permutation of $\mathbf{x}$ can be obtained by permuting the output tokens contained in $H(\mathbf{x})$. We refer to this property as “group permutation symmetry”. If we increase $A$, a greater number of input sequences can be represented as a token permutation of a given input sequence $\mathbf{x}$, thus leading to an increased degree of group permutation symmetry.
In the absence of the said symmetry, the size of the LUT required to store the transformer output would be $(2^{k},n)$, $n$ real values for each of the $2^{k}$ possible input sequences. However, if we do have group permutation symmetry with $A$ tokens of size $\lceil k/A\rceil$, then the size of the LUT is
$\bigg(2^{\lceil k/A\rceil}\binom{2^{\lceil k/A\rceil}+A-2}{2^{\lceil k/A\rceil}-1},\lceil n/A\rceil\bigg).$ (48)
This follows since we have $2^{\lceil k/A\rceil}$ possible input tokens and for each input token, we have an output sequence of length $\lceil n/A\rceil$. Furthermore, the transformer encoder determining the output sequence for the given input token is token permutation in-variant to the other A1A-1 input tokens. If we count the number of distinct A1A-1 input token sequences, ignoring any token permutations, where each token can take $2^{\lceil k/A\rceil}$ distinct values, then that can easily be shown to be $\binom{2^{\lceil k/A\rceil}+A-2}{2^{\lceil k/A\rceil}-1}$.
Consider $k=11$, $n=32$ and $A=4$. Then the reduced size of the LUT is $(960,8)$ which is smaller than $(2^{11},32)$ by a factor of 8.5.
@@ B Learnt PS - Training and Results
We then replace $\sqrt{\alpha_{m}}\mathbf{c}_{m}$ in Algorithm 1 by $\sqrt{\kappa_{\mathbf{\Pi}}}\mathbf{c}_{u,m}$, and the remainder of the training process remains unchanged. If we set $N=1$ and train on $(p,\rho)=(0.9,0)$ (where $\rho\in[0,1]$ is the correlation coefficient between consecutive bits $b_{i-1}$ and $b_{i}$), we will converge to the Learnt PS depicted in Fig. 3. If we perform free-lunch training as described in Section VII-D, the PS learnt (labeled Robust PS) is very similar to Arithmetic PS and is depicted in Fig. 13 for $p=0.9$ and $\rho=\{0,0.5,0.9\}$.
@@
C Bitwise UEP for 1 bit
Consider the vector complex AWGN channel
$\mathbf{y}=\mathbf{x}_{i}+\mathbf{n},\qquad\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{n/2}),$ (50)
so that $\mathbb{E}[\lVert\mathbf{n}\rVert^{2}]=\sigma^{2}n/2$ and equivalently $\Re\{\mathbf{n}\},\Im\{\mathbf{n}\}\sim\mathcal{N}(0,\frac{\sigma^{2}}{2}\mathbf{I}_{n/2})$. Assume binary signaling for the HARQ-ACK bit with repetition coding:
$\mathbf{x}_{1}=+\sqrt{P_{1}}\mathbf{1}_{n/2}\ \text{(ACK)},\qquad\mathbf{x}_{0}=-\sqrt{P_{0}}\mathbf{1}_{n/2}\ \text{(NACK)}.$ (51)
The conditional pdf of $\mathbf{y}$ given $\mathbf{x}_{i}$ is
$p(\mathbf{y}\mid\mathbf{x}_{i})=\frac{1}{\pi^{n/2}\sigma^{2}}\exp\!\left(-\frac{\lVert\mathbf{y}-\mathbf{x}_{i}\rVert^{2}}{\sigma^{2}}\right).$ (52)
A likelihood-ratio test of the form
$\log\frac{p(\mathbf{y}\mid\mathbf{x}_{0})}{p(\mathbf{y}\mid\mathbf{x}_{1})}\geq\log\frac{\beta}{1-\beta}$ (53)
is equivalent to
$\lVert\mathbf{y}-\mathbf{x}_{1}\rVert^{2}-\lVert\mathbf{y}-\mathbf{x}_{0}\rVert^{2}\geq\sigma^{2}\log\frac{\beta}{1-\beta}.$ (54)
Expanding the difference of squared distances yields
$\lVert\mathbf{y}-\mathbf{x}_{1}\rVert^{2}-\lVert\mathbf{y}-\mathbf{x}_{0}\rVert^{2}$ $=\big(\lVert\mathbf{y}\rVert^{2}-2\Re\{\mathbf{x}_{1}^{H}\mathbf{y}\}+\lVert\mathbf{x}_{1}\rVert^{2}\big)-\big(\lVert\mathbf{y}\rVert^{2}-2\Re\{\mathbf{x}_{0}^{H}y\}+\lVert\mathbf{x}_{0}\rVert^{2}\big)$ (55)
$=-2\Re\{(\mathbf{x}_{1}-\mathbf{x}_{0})^{*}y\}+\big(\lVert\mathbf{x}_{1}\rVert^{2}-\lVert\mathbf{x}_{0}\rVert^{2}\big).$
Let $\mathbf{d}\triangleq\mathbf{x}_{1}-\mathbf{x}_{0}$. Substituting (55) into (54) gives the linear test
$\Re\{\mathbf{d}^{H}\mathbf{y}\}\leq\tau,\qquad\tau\triangleq\frac{\lVert\mathbf{x}_{1}\rVert^{2}-\lVert\mathbf{x}_{0}\rVert^{2}-\sigma^{2}\log\frac{\beta}{1-\beta}}{2}.$ (56)
For $\mathbf{x}_{1}=+\sqrt{P_{1}}\mathbf{1}_{n/2}$ and $\mathbf{x}_{0}=-\sqrt{P_{0}}\mathbf{1}_{n/2}$, we have $\lVert\mathbf{x}_{1}\rVert^{2}-\lVert\mathbf{x}_{0}\rVert^{2}=n/2(P_{1}-P_{0})$, and hence $\Re\{\mathbf{d}^{H}\mathbf{y}\}=(\sqrt{P_{1}}+\sqrt{P_{0}})\Re\{\mathbf{1}^{H}\mathbf{y}\}$. Therefore (3) reduces to $\Re\{\mathbf{1}^{H}\mathbf{y}\}\leq\gamma$, where
$\gamma\triangleq\frac{n/2(P_{1}-P_{0})-\sigma^{2}\log\frac{\beta}{1-\beta}}{2(\sqrt{P_{1}}+\sqrt{P_{0}})}.$ (57)
Equivalently, decide ACK iff $\Re\{\mathbf{1}^{H}\mathbf{y}\}>\gamma$. Under NACK($\mathbf{x}_{0}=-\sqrt{P_{0}}\mathbf{1}_{n/2}$), $\Re\{y\}=-\sqrt{P_{0}}\mathbf{1}_{n/2}+\Re\{\mathbf{n}\}$, and thus the NACK error rate is given by
$P(\mathrm{ACK}\mid\mathrm{NACK})=P\bigg(\Re\{\mathbf{1}^{H}\mathbf{y}\}>\gamma\mid\mathbf{x}_{0}\bigg)=Q\!\left(\frac{\gamma+(n/2)\sqrt{P_{0}}}{(n/2)\sigma}\right).$ (58)
Given a target NACK error rate $\delta_{0}$, $\beta$ can be computed as a function of $\sigma^{2}$ and $\delta_{0}$.
D Impact of varying $\beta$ on $\mathrm{SNR}_{\mathrm{th}}$
In Section V-A, we presented an empirical rule for setting $\beta$ as per (15). Fig. 14 illustrates the impact of varying $\beta$ for the NR code with $(k,n)=(11,32)$ in an AWGN channel. If $\beta=0.5\implies\beta/(1-\beta)=1$, then both ACK and NACK are equally protected, hence their BER curves coincide in Fig. 14a. As we lower $\beta$ to 0.25, NACK begins to achieve a lower BER than ACK in Fig. 14b. However, we observe that the SNR required to achieve a 0.1% NACK BER is higher than that required to achieve 1% ACK BER. Finally, at $\beta=1/11$ in Fig. 14c, we observe that the two aforementioned SNRs coincide, which implies that this value of $\beta$ minimizes $\mathrm{SNR}_{\mathrm{th}}$. Furthermore, we observe that $\mathrm{BER}_{\mathrm{NACK}}/\mathrm{BER}_{\mathrm{ACK}}=0.1$ for a band of SNRs around $\mathrm{SNR}_{\mathrm{th}}$. Similarly, we observe that the BER ratio is $1/3$ and 1 in Fig. 14b and 14a respectively. We observed this to be true for other values of $p$ as well, hence we arrived at the rule in (15).
@@ D Impact of varying $\beta$ on $\mathrm{SNR}_{\mathrm{th}}$
E Derivation of (25)
Note that $\mathbf{\Sigma}_{11}=g^{2}\mathbb{E}[\mathbf{h}_{d,l}\mathbf{h}^{H}_{d,l}]$ ($l$ is the receiver antenna index), from which the formula for $\mathbf{\Sigma}_{11}$ in (25) follows. To compute $\mathbf{\Sigma}_{12}$, we have
$\mathbf{\Sigma}_{12}$ $=g\mathbb{E}[\mathbf{h}_{d,l}\mathbf{y}^{H}_{p,l}]$ (59)
$=g\mathbb{E}[\mathbf{h}_{d,l}(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})^{H}]$
$=g^{2}\sqrt{\beta_{m}}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{p}},$
@@ E Derivation of (25)
$=\mathbb{E}[(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})(g\sqrt{\beta_{m}}\mathbf{h}_{p,l}\odot\mathbf{1}_{N_{p}}+\mathbf{n}_{p,l})^{H}]$
$=g^{2}\beta_{m}\mathbf{R}_{\mathbf{h}_{p}\mathbf{h}_{p}}+\sigma^{2}\mathbf{I}_{N_{p}}.$
F Derivation of (36)
@@
G Impact of varying ChEST SNR on RMMSE ChEST